OpenStudy (anonymous):

cscӨcotӨ/secӨ= a)1 b) csc^Ө c)cot^2Ө d)sec^2Ө

OpenStudy (anonymous):

what is this? cscӨ???? = cosecӨ

OpenStudy (anonymous):

c)cot^2Ө

OpenStudy (anonymous):

\[(\cos \theta)(\cos \theta) /(\sin \theta)(\sin \theta)\]

OpenStudy (anonymous):

\[\cot \theta = \cos \theta /\sin \theta\] \[cosec \theta = 1/\sin \theta \] \[1/\sec \theta = \cos \theta \] Put this in the question: You will reach your answer...

OpenStudy (anonymous):

thanks again! that's all for tonight

OpenStudy (anonymous):

so \[\cos ^{2}\theta/\sin ^{2}\theta=\cot ^{2}\theta \]